OpenStudy (anonymous):

I keep on re-doing this problem and getting different answers. Help? Evaluate the expression: x^2-81-px+40 given x=2, p=5. So, 2^2-81-(5)(2)+40 My most recent answers are -27 & -48, but I've also gotten 164, -46, 106, and -124.

OpenStudy (anonymous):

..... 2^2-81-(5)(2)+40 4 -81 -10+40 44-91 -47

OpenStudy (anonymous):

E none of the above

OpenStudy (anonymous):

my suggestion to you, write it down and take it step by step and when in doubt, grab a calculator if you keep making a mistake when adding and subtracting to check your answer as i am assuming that is where you made your mistake on a side note, do not rely on a calculator

OpenStudy (anonymous):

wow my grammar was off there too many thoughts jumbled up in my head use a calculator to check your answer use it as a last resort

OpenStudy (anonymous):

Ok, so I was able to follow along and actually got -47, but you lost me on how you were able to get 44-91. The way I ended up doing it was: 4-81-10+40 -77-10+40 -87+40 = -47 Can you tell me how you were able to get 44-91?

OpenStudy (anonymous):

using the commutative property of addition 4-81-10+40 = 4+40 - 81 - 10 (4+40) (- 81 - 10) 44 - 91

OpenStudy (anonymous):

I had absolutely no idea you could do that. Wow. Okay, I think I get everything now. Thanks a lot!

OpenStudy (anonymous):

commutative property of addition means you can add and subtract in any order

OpenStudy (anonymous):

a+ b = b+ a

OpenStudy (anonymous):

unless there is a parenthesis seperating it in which case you would need to follow order of operation otherwise known as PEMDAS

OpenStudy (anonymous):

Oh, I see. Seriously, thanks so much. I'm going to have to brush up on the properties of addition. Thanks a lot for the tip about parentheses, too. I seriously had no idea you could solve this type of problem that way.

OpenStudy (anonymous):

definitely look up order of operations

OpenStudy (anonymous):

Will do. Once again, thanks a lot. You were extremely helpful.

OpenStudy (anonymous):

no problem